# A rigidity theorem for complex Kleinian groups

### v1

@@ Abstract.
Canary was partially supported by grant DMS-2304636 from the National Science Foundation. Zhang was partially supported by the NUS-MOE grants A-8001950-00-$0$ and A-8000458-00-$0$. Zimmer was partially supported by a Sloan research fellowship and grants DMS-2105580 and DMS-2452068 from the National Science Foundation.
@@ 1. Introduction
Since the property of being (dk)(d-k)-hyperconvex is stable (see [30, Prop. 6.2]), one may view our result as a global analogue of the local rigidity result of Raghunathan [31] (see also Menal–Ferrer and Porti [26, 29]) who proved that if $\Gamma$ is a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, then $\tau|_{\Gamma}$ is locally rigid in the character variety ${\bf X}(\Gamma,\mathsf{PSL}(d,\mathbb{C}))$ if $\tau:\mathsf{PSL}(2,\mathbb{C})\to\mathsf{PSL}(d,\mathbb{C})$ is an irreducible representation.
@@
Farre–Pozzetti–Viaggi [19, Thm. B] showed that if $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex and $\Gamma$ is not virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, then the limit set $\Lambda_{k}(\Gamma)$ of $\Gamma$ has 2-dimensional Hausdorff measure zero. Theorem 1.1 follows immediately from their result and Theorem 1.4.
@@
One key input in the proof of Theorem 1.2 is the fact, which follows from work of Pozzetti–Sambarino–Wienhard [30] (see also Zhang–Zimmer [37]), that if $\Gamma$ is (dk)(d-k)-hyperconvex and $\partial\Gamma$ is a two-sphere, then $\Lambda_{k}(\Gamma)$ is a $\mathcal{C}^{1}$-smooth submanifold. We then observe that $\Lambda_{k}(\Gamma)$ is a complex submanifold of $\mathrm{Gr}_{k}(\mathbb{C}^{d})$, and hence, by Chow’s theorem [14], a complex algebraic subvariety of $\mathrm{Gr}_{k}(\mathbb{C}^{d})$. This implies that the real Zariski closure $\mathsf{G}$ of $\Gamma$ acts by biholomorphisms on $\Lambda_{k}(\Gamma)$. Then, since $\Lambda_{k}(\Gamma)$ is a two-sphere and hence biholomorphic to $\mathbb{CP}^{1}$, we obtain a surjective homomorphism from $\mathsf{G}$ to $\mathsf{PSL}(2,\mathbb{C})$, the biholomorphism group of $\mathbb{CP}^{1}$.
In Section 4 we establish a sufficient condition for the real Zariski closure of an Anosov group to be simple. Using this, we can prove that the homomorphism $\mathsf{G}\rightarrow\mathsf{PSL}(2,\mathbb{C})$ in the proof of Theorem 1.2 is indeed an isomorphism.
@@ 2.2. Flag manifolds and the Iwasawa decomposition
If $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ satisfies $\alpha_{k}(\kappa(g))>0$ for all $k\in\theta$, then we may define $U_{\theta}(g)\in\operatorname{\mathcal{F}}_{\theta}$ by $U_{\theta}(g)^{k}:=U_{k}(g)$ for all $k\in\theta$. The following is standard (see for example, [12, Prop. 2.3]).
@@ 2.3. Divergent, transverse and Anosov subgroups
One can verify that $\Lambda_{\theta}(\Gamma)$ is a $\Gamma$-invariant subset of $\operatorname{\mathcal{F}}_{\theta}$, see [11, Prop. 3.1]. If $\theta$ is symmetric, i.e. $\theta=\iota(\theta)$, a $\mathsf{P}_{\theta}$-divergent subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is called $\mathsf{P}_{\theta}$-transverse if $\Lambda_{\theta}(\Gamma)\subset\operatorname{\mathcal{F}}_{\theta}$ is a transverse subset, i.e. any pair of distinct flags in $\Lambda_{\theta}(\Gamma)$ are transverse.
@@
Using Proposition 2.1, one can verify that if $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-transverse, then it acts on $\Lambda_{\theta}(\Gamma)$ as a convergence group, see [11, Prop. 3.3] or [23, Sec. 5.2]. We say that a $\mathsf{P}_{\theta}$-transverse subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-Anosov if every point in $\Lambda_{\theta}(\Gamma)$ is conical. Equivalently, a $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-Anosov if and only if $\Gamma$ is word hyperbolic and there exists a $\Gamma$-equivariant homeomorphism between $\Lambda_{\theta}(\Gamma)$ and the Gromov boundary $\partial\Gamma$ of $\Gamma$.
@@
The following is a characterization of Anosov representations due to Kapovich–Leeb–Porti [24, Thm. 1.5] and Bochi–Potrie–Sambarino [6, Prop. 4.6 and 4.9].
@@ 2.4. Patterson–Sullivan measures
for $\mu$-almost every $x\in\Lambda_{\theta}(\Gamma)$. By [12, Prop. 3.2], such measures exist.
@@
for every $k\in\{2,\dots,d-2\}$. In the remainder of the paper, when discussing Patterson–Sullivan measures, we will largely specialize to the case when $\Gamma$ is $\mathsf{P}_{\theta_{k}}$-Anosov and $\mu$ is a $\alpha_{k}$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{k}(\Gamma)$. (Notice that $\alpha_{1}=2\omega_{1}-\omega_{2}$, $\alpha_{d-1}=2\omega_{d-1}-\omega_{d-2}$ and $\alpha_{k}=2\omega_{k}-\omega_{k-1}-\omega_{k+1}$ for every $k\in\{2,\dots,d-2\}$, so we may view $\alpha_{k}$ as an element in $\mathfrak{a}_{\theta_{k}}^{*}$ via the usual identification of $\mathfrak{a}_{\theta_{k}}^{*}$ as a subspace of $\mathfrak{a}^{*}$.) Since $\Gamma$ is $\mathsf{P}_{\theta_{k}}$-Anosov and $\alpha_{k}$ is positive on $\mathfrak{a}_{\theta_{k}}^{+}$, it follows from Sambarino [32, Cor. 5.4.3] that $\delta^{\alpha_{k}}(\Gamma)<+\infty$, and [32, Thm. A] that $\mu$ is unique.
@@ 2.5. Hyperconvex groups
With $\theta_{k}\subset\Delta$ as in the previous subsection, a subgroup $\Gamma\subset\mathsf{PSL}(d,\mathbb{K})$ is (dk)(d-k)-hyperconvex if it is $\mathsf{P}_{\theta_{k}}$-Anosov, and for all mutually distinct triples $x,y,z\in\Lambda_{\theta}(\Gamma)$ we have
@@
Theorem 2.4 ([19, Thm. A]).
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is $k$-hyperconvex for some $k\in\{1,\dots,d-1\}$, then $\Gamma$ is virtually isomorphic to a convex cocompact subgroup of $\mathsf{SL}(2,\mathbb{C})$.
@@
If $k\in\{2,\dots,d-1\}$, then there exist $d_{k}$, a representation $\eta_{k}:\mathsf{PSL}(d,\operatorname{\mathbb{K}})\to\mathsf{PSL}(d_{k},\operatorname{\mathbb{K}})$ and a smooth embedding $E_{k}:\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\to\mathbb{\operatorname{\mathbb{P}}}(\operatorname{\mathbb{K}}^{d_{k}})$, which is holomorphic when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$, such that if $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is (dk)(d-k)-hyperconvex subgroup then the following hold.
@@
    $\Lambda_{1}(\eta_{k}(\Gamma))=E_{k}(\Lambda_{k}(\Gamma)).$
@@
    If $\mu$ is the $\alpha_{k}$-Patterson-Sullivan measure on $\Lambda_{k}(\Gamma)$ for $\Gamma$ of dimension $\delta$, then $(E_{k})_{*}(\mu)$ is the $\alpha_{1}$-Patterson-Sullivan measure on $\Lambda_{1}(\eta(\Gamma))$ for $\eta_{k}(\Gamma)$ of dimension $\delta$.
@@ 3. Ahlfors regularity
where $B(x,t)$ is the ball centered at $x$ of radius $t$. In this section we prove that if $\Gamma$ is (dk)(d-k)-hyperconvex, then its $\alpha_{k}$-Patterson–Sulivan measure is Ahlfors regular on $\Lambda_{k}(\Gamma)$.
@@
Suppose that $\delta^{\alpha_{k}}(\Gamma)=2$. By Theorem 3.1, the $2$-dimensional Hausdorff measure of $\Lambda_{k}(\Gamma)$ is positive, so Theorem 2.5 implies that $\partial\Gamma$ is a $2$-sphere. Then by Theorem 2.4, $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.
@@ 3.1. Transverse representations and shadows
If $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a $\mathsf{P}_{1,d-1}$-transverse subgroup, then by making a choice of $d_{0}$, $\Omega$, $\Gamma_{0}$, and $\rho$ as given by Theorem 3.3 (with $k=1$), we may define shadows in $\Lambda_{1}(\Gamma)$ as the images of the shadows in $\Lambda_{\Omega}(\Gamma_{0})$ under $\xi^{1}$, where $\xi=(\xi^{1},\xi^{d-1})$ is the limit map of $\rho$. In our previous work [12, Prop. 7.1], we proved a shadow lemma for these shadows, i.e. we estimated the $\phi$-Patterson–Sullivan measures of the shadows in $\Lambda_{1}(\Gamma)$. Also, assuming an analog of the (d1)(d-1)-hyperconvexity assumption for transverse representations, we estimated the in-radii and out-radii of these shadows [11, Thm. 5.1 and 7.1]. We will now state a specialization of both of these results to the setting where $\Gamma$ is (d1)(d-1)-hyperconvex and $\phi=\alpha_{1}$.
Theorem 3.4 ([12, Prop. 7.1], [11, Thm. 5.1 and 7.1]).
@@
4. Simplicity of the Zariski closure
Notice that $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is the identity component of $\mathsf{PGL}(d,\operatorname{\mathbb{K}})$, which is a real algebraic group. Hence, given a discrete subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, its real Zariski closure in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is the intersection of $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ with the real Zariski closure of $\Gamma$ in $\mathsf{PGL}(d,\operatorname{\mathbb{K}})$. We also say that a discrete subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is strongly irreducible if every finite index subgroup of $\Gamma$ is irreducible. Notice that if $\operatorname{\mathsf{G}}^{0}$ is the identity component of the real Zariski closure of a discrete subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, then $\Gamma^{0}:=\Gamma\cap\operatorname{\mathsf{G}}^{0}$ is a finite index subgroup of $\Gamma$. Thus, if $\Gamma$ is strongly irreducible, then $\Gamma^{0}$, and hence $\operatorname{\mathsf{G}}^{0}$, is irreducible.
In this section we establish a sufficient condition for the real Zariski closure of a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup to be simple.

### v2

@@ Abstract.
Canary was partially supported by grant DMS-2304636 from the National Science Foundation. Zhang was partially supported by the NUS-MOE grants A-8001950-00-00 and A-8000458-00-00. Zimmer was partially supported by a Sloan research fellowship and grants DMS-2105580 and DMS-2452068 from the National Science Foundation.
@@ 1. Introduction
Since the property of being (dk)(d-k)-hyperconvex is stable (see [27, Prop. 6.2]), one may view our result as a global analogue of the local rigidity result of Raghunathan [28] (see also Menal–Ferrer and Porti [24, 26]) who proved that if $\Gamma$ is a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, then $\tau|_{\Gamma}$ is locally rigid in the character variety ${\bf X}(\Gamma,\mathsf{PSL}(d,\mathbb{C}))$ if $\tau:\mathsf{PSL}(2,\mathbb{C})\to\mathsf{PSL}(d,\mathbb{C})$ is an irreducible representation.
@@
Farre–Pozzetti–Viaggi [18, Thm. B] showed that if $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex and $\Gamma$ is not virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, then the limit set $\Lambda_{k}(\Gamma)$ of $\Gamma$ has 2-dimensional Hausdorff measure zero. Theorem 1.1 follows immediately from their result and Theorem 1.4.
@@
One key input in the proof of Theorem 1.2 is the fact, which follows from work of Pozzetti–Sambarino–Wienhard [27] (see also Zhang–Zimmer [35]), that if $\Gamma$ is (dk)(d-k)-hyperconvex and $\partial\Gamma$ is a two-sphere, then the limit set $\Lambda_{k-1,k,k+1}$ of $\Gamma$ in the partial flag variety $\mathcal{F}_{k-1,k,k+1}$ is a $\mathcal{C}^{1}$-smooth submanifold. (Elements of $\mathcal{F}_{k-1,k,k+1}$ are partial flags with subspaces in dimensions k1k-1, $k$ and $k+1$). We then observe that $\Lambda_{k-1,k,k+1}(\Gamma)$ is a complex submanifold of $\mathrm{Gr}_{k}(\mathbb{C}^{d})$, and hence, by Chow’s theorem [13], a complex algebraic subvariety of $\mathcal{F}_{k-1,k,k+1}$. This implies that the real Zariski closure $\mathsf{G}$ of $\Gamma$ acts by biholomorphisms on $\Lambda_{k-1,k,k+1}(\Gamma)$. Then, since $\Lambda_{k-1,k,k+1}(\Gamma)$ is a two-sphere and hence biholomorphic to $\mathbb{CP}^{1}$, we obtain a surjective homomorphism from the identify component $\mathsf{G}^{0}$ of $\mathsf{G}$ to $\mathsf{PSL}(2,\mathbb{C})$, the biholomorphism group of $\mathbb{CP}^{1}$. We use Lie theoretic arguments to prove that $\mathsf{G}$ is simple and then geometric arguments to conclude that $\mathsf{G}$ is isomorphic to $\mathsf{PSL}(2,\mathbb{C})$.
@@ 2.2. Flag manifolds and the Iwasawa decomposition
If $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ satisfies $\alpha_{k}(\kappa(g))>0$ for all $k\in\theta$, then we may define $U_{\theta}(g)\in\operatorname{\mathcal{F}}_{\theta}$ by $U_{\theta}(g)^{k}:=U_{k}(g)$ for all $k\in\theta$. The following is standard (see for example, [11, Prop. 2.3]).
@@ 2.3. Divergent, transverse and Anosov subgroups
One can verify that $\Lambda_{\theta}(\Gamma)$ is a $\Gamma$-invariant subset of $\operatorname{\mathcal{F}}_{\theta}$, see [10, Prop. 3.1]. If $\theta$ is symmetric, i.e. $\theta=\iota(\theta)$, a $\mathsf{P}_{\theta}$-divergent subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is called $\mathsf{P}_{\theta}$-transverse if $\Lambda_{\theta}(\Gamma)\subset\operatorname{\mathcal{F}}_{\theta}$ is a transverse subset, i.e. any pair of distinct flags in $\Lambda_{\theta}(\Gamma)$ are transverse.
@@
Using Proposition 2.1, one can verify that if $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-transverse, then it acts on $\Lambda_{\theta}(\Gamma)$ as a convergence group, see [10, Prop. 3.3] or [21, Sec. 5.2]. We say that a $\mathsf{P}_{\theta}$-transverse subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-Anosov if every point in $\Lambda_{\theta}(\Gamma)$ is conical. Equivalently, a $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-Anosov if and only if $\Gamma$ is word hyperbolic and there exists a $\Gamma$-equivariant homeomorphism between $\Lambda_{\theta}(\Gamma)$ and the Gromov boundary $\partial\Gamma$ of $\Gamma$.
@@
The following is a characterization of Anosov representations due to Kapovich–Leeb–Porti [22, Thm. 1.5] and Bochi–Potrie–Sambarino [6, Prop. 4.6 and 4.9].
@@ 2.4. Patterson–Sullivan measures
for $\mu$-almost every $x\in\Lambda_{\theta}(\Gamma)$. By [11, Prop. 3.2], such measures exist.
@@
In the remainder of the paper, when discussing Patterson–Sullivan measures, we will largely specialize to the case when $\Gamma$ is $\mathsf{P}_{\theta_{k}}$-Anosov and $\mu$ is a $\alpha_{k}$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{k}(\Gamma)$. (Notice that $\alpha_{1}=2\omega_{1}-\omega_{2}$, $\alpha_{d-1}=2\omega_{d-1}-\omega_{d-2}$ and $\alpha_{k}=2\omega_{k}-\omega_{k-1}-\omega_{k+1}$ for every $k\in\{2,\dots,d-2\}$, so we may view $\alpha_{k}$ as an element in $\mathfrak{a}_{\theta_{k}}^{*}$ via the usual identification of $\mathfrak{a}_{\theta_{k}}^{*}$ as a subspace of $\mathfrak{a}^{*}$.) Since $\Gamma$ is $\mathsf{P}_{\theta_{k}}$-Anosov and $\alpha_{k}$ is positive on $\mathfrak{a}_{\theta_{k}}^{+}$, it follows from Sambarino [29, Cor. 5.4.3] that $\delta^{\alpha_{k}}(\Gamma)<+\infty$, and [29, Thm. A] that $\mu$ is unique.
@@ 2.5. Hyperconvex groups
Let $k\in\{1,\dots,d-1\}$. With $\theta_{k}\subset\Delta$ as in the previous subsection, a subgroup $\Gamma\subset\mathsf{PSL}(d,\mathbb{K})$ is (dk)(d-k)-hyperconvex if it is $\mathsf{P}_{\theta_{k}}$-Anosov, and for all mutually distinct triples $x,y,z\in\Lambda_{\theta}(\Gamma)$ we have
@@
Theorem 2.4 ([18, Thm. A]).
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex for some $k\in\{1,\dots,d-1\}$, then $\Gamma$ is virtually isomorphic to a convex cocompact subgroup of $\mathsf{SL}(2,\mathbb{C})$.
@@
If $k\in\{2,\dots,d-1\}$, then there exist $d_{k}$, a representation $\eta_{k}:\mathsf{PSL}(d,\operatorname{\mathbb{K}})\to\mathsf{PSL}(d_{k},\operatorname{\mathbb{K}})$ and a smooth embedding $E_{k}:\operatorname{\mathcal{F}}_{k-1,k,k+1}(\operatorname{\mathbb{K}}^{d})\to\mathbb{\operatorname{\mathcal{F}}}_{1,2}(\operatorname{\mathbb{K}}^{d_{k}})$, which is holomorphic when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$, such that if $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is (dk)(d-k)-hyperconvex subgroup then the following hold.
@@
    $\Lambda_{1,2}(\eta_{k}(\Gamma))=E_{k}(\Lambda_{k-1,k,k+1}(\Gamma)).$
@@
    $E_{k}$ descends to a smooth embedding $E_{k}^{1}:\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\rightarrow\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{k}})$.
  4. (4)
    If $\mu$ is the $\alpha_{k}$-Patterson-Sullivan measure on $\Lambda_{k}(\Gamma)$ for $\Gamma$ of dimension $\delta$, then $(E_{k}^{1})_{*}(\mu)$ is the $\alpha_{1}$-Patterson-Sullivan measure on $\Lambda_{1}(\eta(\Gamma))$ for $\eta_{k}(\Gamma)$ of dimension $\delta$.
@@ 3. Ahlfors regularity
where $B(x,t)$ is the ball centered at $x$ of radius $t$. Let $k\in\{1,\dots,d-1\}$. In this section we prove that if $\Gamma$ is (dk)(d-k)-hyperconvex, then its $\alpha_{k}$-Patterson–Sulivan measure is Ahlfors regular on $\Lambda_{k}(\Gamma)$.
@@
Suppose that $\delta^{\alpha_{k}}(\Gamma)=2$. By Theorem 3.1, the $2$-dimensional Hausdorff measure of $\Lambda_{k}(\Gamma)$ is positive, so Theorem 2.5 implies that $\partial\Gamma$ is a 22-sphere. Then by Theorem 2.4, $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.
@@ 3.1. Transverse representations and shadows
If $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a $\mathsf{P}_{1,d-1}$-transverse subgroup, then by making a choice of $d_{0}$, $\Omega$, $\Gamma_{0}$, and $\rho$ as given by Theorem 3.3 (with $k=1$), we may define shadows in $\Lambda_{1}(\Gamma)$ as the images of the shadows in $\Lambda_{\Omega}(\Gamma_{0})$ under $\xi^{1}$, where $\xi=(\xi^{1},\xi^{d-1})$ is the limit map of $\rho$. In our previous work [11, Prop. 7.1], we proved a shadow lemma for these shadows, i.e. we estimated the $\phi$-Patterson–Sullivan measures of the shadows in $\Lambda_{1}(\Gamma)$. Also, assuming an analog of the (d1)(d-1)-hyperconvexity assumption for transverse representations, we estimated the in-radii and out-radii of these shadows [10, Thm. 5.1 and 7.1]. We will now state a specialization of both of these results to the setting where $\Gamma$ is (d1)(d-1)-hyperconvex and $\phi=\alpha_{1}$.
Theorem 3.4 ([11, Prop. 7.1], [10, Thm. 5.1 and 7.1]).
@@
4. Rigidity
In this section, we complete the proof of Theorem 1.2 by establishing our rigidity result in the lattice case. Recall that a discrete subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is strongly irreducible if every finite index subgroup of $\Gamma$ is irreducible. As before, let $k\in\{1,\dots,d-1\}$.